OpenStudy (anonymous):

I really need help solving this equation, Im confused of the steps 5v-10/21 divided by v-2/77

OpenStudy (watchmath):

Dividing by \(\frac{v-2}{77}\) is the same as multiplying by \(\frac{77}{v-2}\). So we get \[\frac{5(v-2)}{21}\cdot\frac{77}{v-2}=\frac{5\cdot 77}{21}=\frac{35}{3}\]

OpenStudy (watchmath):

Sorry the final answer should be \(\frac{55}{3}\).

OpenStudy (anonymous):

thank u